OpenStudy (anonymous):

The perimeter of a rectangle is the sum of all of its sides. Write a mathematical expression for the perimeter of a rectangle with a length of 4 and a width of w. Simplify the expression if possible.(

OpenStudy (anonymous):

8+2w

OpenStudy (anonymous):

P=2l+2w P=2*4+2*w=8+2w